If sinx = (a - 1)/x and cosx = (1 - a)/x. Find tanx
Answer by stanbon(75887)

If sinx = (a - 1)/x and cosx = (1 - a)/x.
Find tanx
-----
tan(x) = sin(x)/cos(x) = [(a-1)/x] / [(1-a)/x]
=====
= [(a-1)/x] / [-(a-1)/x]
-----
= -1
================
Cheers,
Stan H.
================